OpenStudy (anonymous):

Two single-digit random numbers (0 to 9 inclusive) are selected independently. Find the probability that their sum is 7. (the answer is 2/25, why?)

OpenStudy (anonymous):

For the first number: 9 and 8 don't work, so you must pick one of the other 8 =8/10 For the second number: Again, 9 and 8 don't work. Also, there is only ONE number that will result in a sum of 7 =1/10 1/8 * 1/10= 2/25

OpenStudy (zarkon):

MathUser has a small typo last line should read \[\frac{8}{10}\times\frac{1}{10}=\frac{8}{100}=\frac{2}{25}\]